# Rethinking Causal Discovery Through the Lens of Exchangeability

### v1

@@ 2.3 Previous Work Intersecting Causality and Exchangeability
Currently, the concepts of causality and exchangeability overlap in two principal domains: the well-established field of causal inference, which fundamentally relies on exchangeability, and the recent work that generalizes de Finetti’s theorem to the causal case and introduces a novel algorithm for causal discovery (Guo et al., 2024).
In the experimental setting, often the goal is to perform cause-effect identification. This means inferring whether a variable has a causal effect on another, while knowing the reverse is not possible (for example, if one variable corresponds to an event that precedes the other). This is often assessed using randomized trials, which can be viewed through the potential outcomes framework (Rubin, 2005). Suppose one wants to assess if a certain treatment ($T$) is effective against a specific disease (outcome $Y$). To do so, a common approach is to gather a group of patients, of which half takes the treatment ($T=1$) and another half does not ($T=0$); the latter is usually called the control group. From this data, the goal is to try to estimate whether the treatment was effective or not. This is given by the average treatment effect (ATE) defined as (Saarela et al., 2023)
$\text{ATE}=\mathbb{E}_{D}[Y(1)-Y(0)]=\mathbb{E}_{D}[Y(1)]-\mathbb{E}_{D}[Y(0)].$ (4)
where $\mathbb{E}_{D}$ is the expected value of each variable over the entire distribution, and $Y(1)$ and $Y(0)$ are two random variables representing the outcome for $T=1$ and $T=0$ respectively.
The main challenge lies in computing this expression since, for each patient, it is either observed $Y(1)$ or $Y(0)$, never both. Additionally, it obviously makes no sense to consider each patient as a different i.i.d. variable of the same process, given the relevant differences between patients, such as age and health condition.
In order for ATE to be computable, exchangeability between the different patients must be assumed (Rosenbaum and Rubin, 1983). To see why, all that is needed is to manipulate the original exchangeability equality (Equation 2.1). Suppose the indices up to n belong to the control group, and the goal is to generalize the results for the treatment group in the control group. This can be done because
@@
$\iff\mathbb{E}_{\{(t,y)\in D:t=1\}}[Y(1)]=\mathbb{E}_{\{(t,y)\in D:t=0\}}[Y(1)].$ (7)
Therefore, the fact the data is exchangeable implies that $\mathbb{E}_{\{(t,y)\in D:t=1\}}[Y(1)]$ can replace $\mathbb{E}_{D}[Y(1)]$ (and similarly for the control group) (Hernán and Robins, 2020), which allows computing the ATE. More intuitively, the data being exchangeable means that it is fair to generalize the results obtained in one group to the entire population. This is valid since the patients in both groups were drawn from the same latent distribution (Höfler, 2005). In a more basic level, this also justifies the need for the patients to be picked at random, so that their choice is not informative over the underlying latent variables distribution.
Additionally, the converse is also true: if the data is not exchangeable, then it is impossible to compute the ATE, since it is impossible to generalize to the entire population. If this were the case, it would mean there was some difference in the latent variable distribution between both groups (e.g. one group could be younger than the other). This would invalidate extrapolating the results obtained from one group to the other, as this could lead to misleading results. Nevertheless, if one is aware of such imbalances, it is possible to adjust for them by considering them as observed variables (if they are available), and performing the exact same reasoning while conditioning on these observed variables (Lee and Lee, 2022).
Interestingly, within the potential outcomes framework, exchangeability implies that correlation is equivalent to causation (Rosenbaum and Rubin, 1983; Hernán and Robins, 2020). This is because under the exchangeability hypothesis, all latent variables (which could generate confounding) are controlled for by both groups following the same prior.
The connection between exchangeability and causal discovery was made only recently: in work that remains the only joining both fields, Guo et al. adapt de Finetti’s Theorem (see Section 2) to the setting of causal discovery.
@@
Let $\{(X_{n},Y_{n})\}_{n\in\mathbb{N}}$ be an infinite sequence of binary random variable pairs. The sequence satisfies the following properties: 1) It is infinitely exchangeable; 2) $\forall n\in\mathbb{N}:Y_{[n]}\perp\!\!\!\perp X_{n+1}\mid X_{[n]}$, where $n]=\{1,\ldots,n\}$, if and only if there exist two random variables $\theta\in[0,1]$ and $\psi\in[0,1]$, with probability measures $\mu$ and $\nu$, respectively, such that the joint probability can be represented as:
$P(X_{1},Y_{1},\dots,X_{N},Y_{N})=\int_{\Phi}\int_{\Theta}\prod_{n=1}^{N}p(Y_{n}\mid X_{n},\psi)p(X_{n}\mid\theta)\,d\mu(\theta)\,d\nu(\psi).$ (8)
In (8), $d\mu(\theta)$ and $d\nu(\psi)$ can be substituted by $p(\theta)d\theta$ and $p(\psi)d\psi$, respectively, since $\mu$ and $\nu$ will always refer to absolutely continuous densities w.r.t. Lebesgue measure on $\Theta$ and $\Psi$.
@@
Additionally, Guo et al. also proposed an algorithm for causal discovery in the bivariate setting based on the asymmetry between both causal directions present in the second assumption of the Causal de Finetti Theorem. The idea is that, while in the causal direction, $Y_{[n]}\perp\!\!\!\perp X_{n+1}\mid X_{[n]}$ holds, the same is not necessarily true in the opposite direction. To test whether $\forall n\in\mathbb{N}:Y_{[n]}\perp\!\!\!\perp X_{n+1}\mid X_{[n]}$, one would need to have the joint probability distribution over $P(Y_{[n]},X_{n+1},X_{[n]})$. However, the asymmetry expressed above is impossible to test in an exchangeable process since only one realization of each variable is obtained in reality. To go around this issue, the authors constructed datasets that approximate de Finetti’s Theorem in viewing an exchangeable process as a mixture of i.i.d. sequences. To do so, they assume the data comes from N different environments, where each environment, $e$, is defined by its latent parameters, which are drawn according to the prior distribution: $(\theta^{e},\psi^{e})\sim(p(\theta),p(\psi))$. In a sense, this is similar to the earlier concept of interventions, because all environments share the same underlying causal graph over the observed variables $X$ and $Y$, while the specific causal mechanism is defined by $\theta$ and $\psi$. Nevertheless, from a more practical viewpoint, these interventions are much softer between each other (very similar causal relationships between interventions), and many more in number (the author’s algorithm considers 100).
3 Exchangeability in Causal Discovery
This paper aims to recast i.i.d. causal discovery in terms of exchangeable data. We begin by examining the potential consequences of this reformulation and then offer a series of arguments to justify its adoption.
3.1 Implications
Since exchangeable sequences may not be i.i.d., it is important to consider what properties and differences exchangeable sequences hold in general, which are absent in i.i.d. sequences.
The first relevant difference is that, in exchangeable sequences, the different samples may not be independent from each other. Guo et al. proposed a causal discovery algorithm by noticing that a certain conditional independence relationship between different samples only holds in one causal direction (Guo et al., 2024), as explained in Section 2.3. New methods for causal discovery could be developed by identifying and mathematically formalizing asymmetries in data that are induced by the direction of the underlying causal relationship. Additionally, it is also relevant to point out that all causal discovery methods that are valid for the i.i.d. case should also be valid for exchangeable data. This stems from these methods relying upon an even more restrictive assumption, which holds in the causal direction for i.i.d. data; therefore, it is naturally generalizable to exchangeable data. On the other hand, the opposite is not necessarily true. Causal discovery methods developed for exchangeable data might take advantage of asymmetries that disappear when we focus on i.i.d. data (see Section 2.3). Therefore, considering the data to be exchangeable instead of i.i.d. opens the door to new methods that may take advantage of new patterns in the causal relationships, while ensuring the verified methods for i.i.d. data keep their validity.
From a different angle, the de Finetti theorem (see Section 2.3) states that any exchangeable sequence can be written as a continuous mixture of i.i.d. sequences. This mixture is controlled by unobserved latent variables. At first glance, these latent variables might seem to be exactly the same as the exogenous variables in the SCM and FCM models, which are usually interpreted as noise; in fact, mathematically speaking, this is the case. However, noise is traditionally used with a very different connotation from the notion of latent variables. For one, noise tends to be thought of as having a small effect in the data, whilst the existence of a latent variable is generally assumed to be as significant as the observed ones. Secondly, noise is also generally believed to affect the causal function in simple ways (one often considers noise to be additive or multiplicative), whilst a latent variable is considered to introduce significant complexity in this relationship. Therefore, even though there is no strict difference between naming the exogenous variables as noise or latent variables, there are many design disparities that may stem from this difference. Consequently, considering the data to be exchangeable would imply the existence of latent variables, which could motivate causal discovery methods to ponder their existence explicitly and thus take into account the variety of complex effects they might have on the observed joint distribution.
Lastly, it is interesting to note that exchangeability between samples can also be linked with the idea of interventions (see Section 1). One can think of an exchangeable sequence as corresponding to each sample being obtained from a different environment (different intervention). In this scenario, each environment would be defined by the latent variable, which in turn changes the causal relationship between the observed variables.
@@ 3.2 Arguments
3.2.1 Conceptual Argument
@@
The conceptual argument in favor of framing i.i.d. causal discovery in exchangeable data starts by noticing the clear similarities between causal inference and causal discovery. Crucially, they share the same setup and the same ultimate goal. The difference is that in causal inference, it is possible to intervene directly in the system, while causal discovery relies solely on observational data. Since they operate under the same formal assumptions and objectives, it makes sense to require the same foundational properties of the data in both settings. In experimental causal inference, one of the key foundational assumptions is exchangeability, which allows for controlling for external factors (see Section 2.3). It has been widely accepted and thoroughly validated in the scientific community as the criterion that enables extrapolating causal conclusions drawn from experimental data (Saarela et al., 2023).
While experimental causal inference has long been a routine, rigorously validated practice, observational causal discovery remains primarily an exploratory research field. Consequently, observational causal discovery should look towards causal inference in order to extrapolate verified and meaningful assumptions on the data. More specifically, it should resemble causal inference in considering exchangeability to be the key statistical property of observational data.
This perspective contrasts sharply with the traditional assumption that the observational data are independent and identically distributed (i.i.d.), since this would require that the system is completely isolated. In turn, the introduction of exchangeability opens the possibility of considering the existence of external unobserved variables. This precise existence is supported by causal inference posing exchangeability as one of its foundational assumptions. Consequently, a failure to recognize such latent variables will necessarily result in the development of extremely narrow-sighted causal discovery algorithms.
3.2.2 Empirical Argument
The empirical argument for reframing i.i.d. causal discovery using exchangeability draws on recurring patterns in recent causal-discovery research that become clearer under this lens. First, we re-examine the Tübingen dataset and show that most of its examples are better characterized as exchangeable rather than strictly i.i.d. Then, we survey state-of-the-art methods to identify those that implicitly exploit exchangeability, typically by modulating the latent variable at the heart of de Finetti’s theorem.
As mentioned in Section 2.2, the Tübingen dataset collection is the only extensive real-world i.i.d benchmark present in the literature and widely used to assess bivariate causal discovery methods. Consequently, we decided to analyze it closely, in order to understand whether the examples it includes are strictly i.i.d. or whether their samples should be considered exchangeable. From the de Finetti theorem (Section 2), exchangeability implies the existence of latent variables, whilst i.i.d. data only occurs when the system is isolated. Therefore, we checked if there existed any plausible unobserved variable in each example of the dataset (the full classification can be seen in Appendix A). After doing so, we concluded that the vast majority of the datasets in the collection ($81.5\%(74.4\%)$ [^5] The second result corresponds to the weighted average, according to the weights proposed by the authors of the dataset.) should be actually considered as exchangeable, while there was a small percentage containing timeseries ($10.2\%(15.4\%)$) and another that can considered i.i.d. ($8.3\%(10.3\%)$).
Several exchangeable examples can be seen in datasets relating geographical features (either latitude, longitude, or altitude) with meteorological quantities (temperature and precipitation). In this scenario, the hidden geographical features are clearly latent variables. A similar situation happens in the datasets connecting the compressive strength of cement with each of its mixture components, or the fuel economy of cars with each of its many features. Additionally, there are many other examples that have very reasonable latent variables. The sets containing the associations between different features of living beings (humans and oysters in specific) with their age clearly involve latent variables in the genetics of each individual.
On the other hand, there are two sets of examples that tried to eradicate the influence of latent variables. One relates different physical attributes of a rolling ball (such its height), with the initial and final velocity, while the other contains both measurements of the emitted and perceived light by a block of LEDs and a photo receptor.
Finally, some datasets fall in the category of timeseries, such as those relating inside and outside temperature or those that associate temperature with the day of the year. As we know, timeseries are not exchangeable, so these could undermine the conclusions we were trying to reach. However, since causal discovery with i.i.d. data is a slightly separate field of research from in timeseries, we believe these examples are present in this dataset by mistake. In principle, any method that works well on i.i.d. or generally exchangeable data should also perform well on timeseries. In practice, however, explicitly recognizing the sequential nature of timeseries lets us design algorithms that exploit far richer temporal dependencies—and ultimately uncover more nuanced causal relationships. The results of our analysis of the Tübingen dataset are shown in Figure 1.
Figure 1: On the left, we can see the unweighted and weighted distribution of statistical assumptions, while on the right, we have 2 examples of timeseries present in the Tübingen dataset.
[…]
The second implication of reframing i.i.d. causal discovery as exchangeable pertains to the difference in interpreting unobserved variables as latent variables instead of noise (see Subsection 3.1). Therefore, methods that allow the unobserved variables to influence the effect in complex and strong ways, can be seen as assuming the existence of latent variables. This allows them to be interpreted in the light of exchangeability, not exactly satisfying the i.i.d. assumption in this area of causal discovery.
The cleanest example in the state of the art is CGNN (Goudet et al., 2018). This method chooses the causal graph whose generated distribution using a neural network lies closest to the true distribution. Both observed and unobserved variables are considered as inputs to the neural network. Therefore, there is complete symmetry between what is observed and unobserved (the only difference is that the marginal distribution of observed variables is obtained from the data, while that of unobserved variables is assumed according to prior knowledge). Therefore, the structure of the neural network ensures that the unobserved variables can contribute in exactly the same way to the effect as the observed causes.
Another interesting approach to exchangeability is bCQD (Tagasovska et al., 2020), which attempts to perform causal discovery by comparing the complexity of enconding the joint distribution using the two possible causal factorizations [^7] The two possible causal factorizations are: $P(X,Y)=P(Y\mid X)P(X)$ and $P(X,Y)=P(X\mid Y)P(Y)$.. However, and more relevant to the subject of this paper, it models the function as a set of quantile regressions. The underlying rationale can only be understood if we consider the existence of a latent variable, and each different quantile regression is trying to model the system for a fixed value of the latent variable. Therefore, assuming exchangeability also contributes to a clearer understanding of this method.
It is important to mention, however, that while some methods are much better understood by extrapolating to outside the i.i.d. domain, this is not generally the case. There are also a variety of methods whose main assumptions are placed in the nature of the causal function, such as SLOPE (Marx and Vreeken, 2017) and IGCI (Mian et al., 2023). In other methods, the noise nature of the unobserved variables (weak impact and low complexity) are a key part of the algorithm (see RECI (Blöbaum et al., 2018) or LiNGAM (Shimizu et al., 2006)). Nonetheless, this is to be expected given how the causal discovery problem is formulated for i.i.d. data. We believe that the emergence of methods based on exchangeability within this domain, achieving performance comparable to other state-of-the-art algorithms [^8] See Guyon et al. and Tagasovska et al. for a comparative analysis of different methods(Guyon et al., 2019; Tagasovska et al., 2020)., serves as further evidence supporting our central argument. Moreover, we are hopeful that extending i.i.d. causal discovery to the exchangeability framework will open up opportunities for many additional methods to be developed.

### v2

@@ 2.4 Previous Work Intersecting Causality and Exchangeability
Currently, the concepts of causality and exchangeability overlap in two principal domains: the well-established field of causal inference, which fundamentally relies on exchangeability, and the recent work by Guo et al. that generalizes de Finetti’s theorem to the causal case and introduces a novel approach to causal discovery (Guo et al., 2024).
2.4.1 Exchangeability in Average Treatment Effect Estimation
In the experimental setting, the goal is often to infer whether a variable has a causal effect on another, knowing the reverse is impossible (e.g., one variable corresponds to an event that precedes the other) . This is often assessed using randomized trials that can be viewed through the potential outcomes framework (Rubin, 2005). Suppose the goal is to assess the effect of a treatment ($T$) on the outcome of some disease ($Y$). The common approach is to gather a group of patients, of which a randomly selected half is subject to the treatment ($T=1$), and the other half (the control group) is not ($T=0$). The outcome ($Y\in\{0,\,1\}$) is recorded for each patient. From this data, the goal is to estimate the so-called average treatment effect (ATE), defined as
$\text{ATE}=\mathbb{E}_{D}[Y(1)-Y(0)]=\mathbb{E}_{D}[Y(1)]-\mathbb{E}_{D}[Y(0)],$ (11)
where $\mathbb{E}_{D}$ denotes expected value over the entire distribution, and $Y(1)$ and $Y(0)$ are two random variables representing the outcome for $T=1$ and $T=0$ respectively (Holland, 1986). The main challenge lies in estimating this quantity (even in the presence of infinitely many patients) since, for each patient, either $Y(1)$ or $Y(0)$ is observed, never both. Additionally, it makes no sense to consider the patients as i.i.d. samples, given the relevant differences between them, such as age and health condition.
To estimate the ATE, it is sufficient to assume exchangeability (Saarela et al., 2023; Rosenbaum and Rubin, 1983), as reviewed next. Suppose the indices up to $N$ belong to the control group and those from $N+1$ to $2N$ to the treatment group, and that the goal is to generalize the results of the treatment group to the control group. Begin by noticing that exchangeability guarantees the following equality:
@@ 2.4.1 Exchangeability in Average Treatment Effect Estimation
Marginalizing both sides of the equality with respect to the second half of the arguments yields
$P(Y_{1}(1),\ldots,Y_{N}(1))=P(Y_{N+1}(1),\ldots,Y_{2N}(1)),$ (13)
which implies that
$\mathbb{E}_{D}[Y(1)|T=0]=\mathbb{E}_{D}[Y(1)|T=1]=\mathbb{E}_{D}[Y(1)].$ (14)
Therefore, exchangeability implies that $\mathbb{E}[Y(1)|T=1]$ can replace $\mathbb{E}_{D}[Y(1)]$ and similarly for the control group ($\mathbb{E}[Y(0)|T=0]$ can replace $\mathbb{E}_{D}[Y(0)]$), which allows computing the ATE (Hernán and Robins, 2020). More intuitively, exchangeability means that it is fair to generalize the results obtained in one group to the entire population. This is valid since the patients in both groups were drawn from the same latent distribution (Höfler, 2005).
Additionally, the converse is also true: if the data is not exchangeable, then it is impossible to estimate the ATE, since there is no valid argument to generalize to the entire population. For example, if there latent distributions of both groups are different (e.g., one group could be younger than the other). This would invalidate extrapolating the results obtained from one group to the other, which could lead to misleading results. Nevertheless, if one is aware of such imbalances, it is possible to adjust for them by considering them as observed variables (if they are observed), and performing the exact same reasoning while conditioning on these observed variables (Lee and Lee, 2022).
Interestingly, within the potential outcomes framework, exchangeability implies that correlation is equivalent to causation (Rosenbaum and Rubin, 1983; Hernán and Robins, 2020). This is because, under the exchangeability hypothesis, all latent variables (which could generate confounding) are controlled for by both groups following the same prior.
2.4.2 The Causal de Finetti Theorem
@@
$P(X_{\pi(1)},Y_{\pi(1)}\ldots,X_{\pi(N)},X_{\pi(N)})=P(X_{1},Y_{1},\ldots,X_{N},Y_{N}).$ (15)
@@
Let $\{(X_{n},Y_{n})\}_{n\in\mathbb{N}}$ be an infinite sequence of binary random variable pairs. The sequence satisfies the two following properties
  1. 1.
    infinite exchangeability, and
  2. 2.
    $\forall n\in\mathbb{N}:Y_{[n]}\perp\!\!\!\perp X_{n+1}|X_{[n]}$, where $n]=\{1,\ldots,n\}$,
@@
if and only if there exist two random variables $\theta\in[0,1]$ and $\psi\in[0,1]^{2}$, with probability measures $\mu$ and $\nu$, respectively, such that
$P(X_{1},Y_{1},\dots,X_{N},Y_{N})=\int_{\Phi}\int_{\Theta}\prod_{n=1}^{N}p(Y_{n}|X_{n},\psi)p(X_{n}|\theta)\,d\mu(\theta)\,d\nu(\psi).$ (16)
In (16), $d\mu(\theta)$ and $d\nu(\psi)$ can be substituted by $p(\theta)d\theta$ and $p(\psi)d\psi$, respectively, if $\mu$ and $\nu$ are absolutely continuous w.r.t. the Lebesgue measure on $\Theta$ and $\Psi$. Furthermore, as in the original de Finetti theorem, if both $\Theta$ and $\Phi$ are discrete sets, Equation (16) takes the form
$P(X_{1},Y_{1},\dots,X_{N},Y_{N})=\sum_{\phi\in\Phi}\sum_{\theta\in\Theta}\prod_{n=1}^{N}p(Y_{n}|X_{n},\psi)p(X_{n}|\theta)\,p(\theta)\,p(\psi),$ (17)
where $p(\theta)$ and $p(\psi)$ are probability mass functions.
@@
Guo et al. also argue that eschewing stronger i.i.d. constraints for exchangeability allows the presence of asymmetric dependencies between variables. These asymmetries can then be explored to develop new causal discovery methods. Specifically, they propose an algorithm for bivariate causal discovery based on the asymmetry between both causal directions present in the second assumption of the Causal de Finetti Theorem. While in the causal direction, $Y_{[n]}\perp\!\!\!\perp X_{n+1}|X_{[n]}$ holds, the same is not necessarily true in the opposite direction. To test whether $\forall n\in\mathbb{N}:Y_{[n]}\perp\!\!\!\perp X_{n+1}|X_{[n]}$, one would need to have the joint probability distribution over $P(Y_{[n]},X_{n+1},X_{[n]})$. However, the asymmetry expressed above is impossible to test in an exchangeable process, since only one realization of each variable is obtained. To address this issue, the authors constructed datasets that approximate de Finetti’s Theorem in viewing an exchangeable process as a mixture of i.i.d. sequences. Specifically, they assume the data comes from $N$ different environments, where each environment, $e$, is defined by its latent parameters, which are drawn according to the prior distribution: $(\theta^{e},\psi^{e})\sim(p(\theta),p(\psi))$. In a sense, this is similar to the earlier concept of interventions, because all environments share the same underlying causal graph over the observed variables $X$ and $Y$, while the specific causal mechanism is defined by $\theta$ and $\psi$. Nevertheless, from a more practical viewpoint, these interventions are much softer between each other (very similar causal relationships between interventions), and many more in number (the authors’ algorithm considers 100).
3 Exchangeability for Causal Discovery
3.1 Hierachical Model
In this section, with the aim of offering a more complete and flexible description of causal processes, a new hierarchical model is proposed. This model extends the Causal de Finetti representation theorem (Equation (16)) in a similar way that the hierarchical models in Subsection 2.3 extend the original de Finetti theorem (Equation (3)). The proposed formulation takes the terms $p(X_{n}|\theta)$ and $p(Y_{n}|X_{n},\psi)$ as representing two marginal probability distributions, integrated over inner latent variables. With $\eta_{n}$ and $\gamma_{n}$ representing these latent variables, marginalization with respect to them yields
$P(X_{1},Y_{1},\dots,X_{N},Y_{N})$ (18)
$=$ $\int_{\Phi}\int_{\Theta}\prod_{n=1}^{N}\left(\int_{H}p(X_{n}|\eta_{n},\theta)\,p(\eta_{n}|\theta)\,d\eta_{n}\right)\left(\int_{\Gamma}p(Y_{n}|X_{n},\gamma_{n},\psi)\,p(\gamma_{n}|\psi)\,d\gamma_{n}\right)\,p(\theta)p(\psi)\,d\theta\,d\psi,$
where
$\int_{H}p(X_{n}|\eta_{n},\theta)\,p(\eta_{n}|\theta)\,d\eta_{n}=p(X_{n}|\theta),$ (19)
and
$\int_{\Gamma}p(Y_{n}|X_{n},\gamma_{n},\psi)\,p(\gamma_{n}|\psi)\,d\gamma_{n}=p(Y_{n}|X_{n},\psi).$ (20)
As above, the integrals in these three equations are replaced by sums in the case of discrete variables.
This hierarchical model provides more flexibility, while still allowing for efficient sampling. The equations above are purposefully general: the variables ($\eta_{n},\gamma_{n},\theta,\psi$) can be scalars or vectors, continuous, discrete, or mixed, depending on the exact nature and formulation of the problem being addressed. Namely, the variable $\theta$ plays two roles in that it defines the distribution of $\eta_{n}$ and the conditional distribution of $X_{n}$ given $\eta_{n}$. The same holds for $\psi,$ which defines the distribution of $\gamma_{n}$ as well as the conditional distribution of $Y_{n}$, given $X_{n}$ and $\gamma_{n}$ (i.e., the causal mechanism). A graphical representation of this model, alongside the exchangeable model proposed by Guo et al. (Guo et al., 2024), is shown in Figure 2.
[…]
Figure 2: Left: exchangeable model of Guo et al. (Guo et al., 2024). Right: proposed exchangeable hierarchical model. The plate represents N i.i.d. samples, while the variables in white are observed, while those in grey are not.
It is important to note that $p(Y_{n}|X_{n},\psi)$ in Equation (20) can entail two fundamentally different types of interactions. Analogously to the previous example, in the term $p(\gamma_{n}|\psi)$, $\psi$ defines the shape and parameters of the inner latent variable $\gamma_{n}$ (e.g., if $\psi<0$, $p(\gamma_{n}|\psi)=\mathcal{N}(0,1)\,\text{; if}\,\psi>0,p(\gamma_{n}|\psi)=\mathcal{U}(0,1)$). On the other hand, $\psi$ can also select (or inform) over the causal mechanism between $X_{n}$ and $Y_{n}$ (e.g., if $\psi$ is smaller than 0, then the $Y_{n}=X_{n}^{2}$, otherwise $Y_{n}\sim\mathcal{B}(X_{n},\psi^{2})$). These two different and complementary interpretations of Equation (18) allow for the proposal of two factors that should lead causal sequences to be considered exchangeable and not i.i.d.: 1) when the causal mechanism between $X_{n}$ and $Y_{n}$ is unknown, and 2) when there exist latent variables whose distribution is unknown. This stems from the fact that the lack of knowledge over either of these two components can be accurately modeled using an exchangeable sequence and, more precisely, the hierarchical model presented in this section.
@@ 3.2 Arguments
The key contribution of this work is the proposal that i.i.d. causal discovery is more accurately framed as exchangeable, in accordance with the hierarchical model defined in Subsection 3.1. According to this model, it suffices for either the causal mechanism between $X_{n}$ and $Y_{n}$, or the distribution of any existing latent variables, to be unknown to break the i.i.d. assumption. In fact, we believe it is reasonable to argue that this is most often the case in real data. Bivariate real-world examples tend to stem from very complex causal networks, where both mechanisms and latent variables are not easy to identify, often being left as underlying assumptions. In fact, if only considering the existence of latent variables with an unknown distribution, an analysis of the Tübingen collection led to the conclusion that $74.4\%$ of datasets are more accurately described as exchangeable, while $10.3\%$ have no clear latent variables with unknown distribution, and $15.3\%$ are time series (see Appendix A for a full breakdown). These statistics in the Tübingen collection, the only real-world bivariate causal discovery dataset, further consolidate the core claim of this work.
@@
An additional argument stems from the clear similarities between causal inference and causal discovery. Crucially, they share the same setup and the same ultimate goal. The difference is that, in causal inference, it is possible to intervene directly in the system, while causal discovery relies solely on observational data. Since they operate under the same formal assumptions and objectives, it makes sense to require the same foundational properties of the data in both settings. In experimental causal inference, one of the key assumptions is exchangeability (see Section 2.4.1). It enables controlling for latent factors without necessitating the assumption of their absence, which is essential for addressing the complexities of real-world data. As such, the exchangeability assumption in causal inference has been widely accepted and thoroughly validated in the scientific community as the criterion that enables extrapolating causal conclusions drawn from experimental data (Saarela et al., 2023). While experimental causal inference has long been a routine, rigorously validated practice, observational causal discovery remains primarily an exploratory research field. Consequently, observational causal discovery should look towards causal inference in order to extrapolate verified and meaningful assumptions from the data. More specifically, it should resemble causal inference in considering exchangeability to be a key statistical property of observational data.
The last argument will rely on the expressiveness of a synthetic dataset generated in accordance with the hierarchical model defined in Equation (18). By relaxing the stricter i.i.d. assumption, the hierarchical model enables the synthetic generation process to encompass a broader family of causal mechanisms, better reflecting the inherent complexities of real-world data. Having the Tübingen dataset as a baseline, we will first discuss how different methods score more closely (in comparison with the Tübingen) relative to other synthetic datasets (Section 4). Then, in Section 5, we will introduce a new causal discovery method: a neural network trained exclusively on the developed synthetic dataset, and show how it performs similarly to other state-of-the-art methods (Section 5).